# Information Discovery in e-Commerce

### v1

@@ 3.1.2.6 User reviews in e-commerce
Opinion tags refer to a ranked list of tags provided by the e-commerce platform that reflect the characteristics of reviews of an item; see Fig. 3.15. To assist consumers to quickly grasp a large number of reviews about an item, opinion tags are increasingly being applied by e-commerce platforms. Current mechanisms for generating opinion tags rely on either manual labelling or heuristic methods, which is time-consuming and ineffective. li2021abstractive proposed the abstractive opinion tagging task, where systems have to automatically generate a ranked list of opinion tags that are based on, but need not occur in, a given set of user-generated reviews.
The abstractive opinion tagging task comes with three main challenges: (i) the noisy nature of reviews; (ii) the formal nature of opinion tags vs. the colloquial language usage in reviews; and (iii) the need to distinguish between different items with very similar aspects. To address these challenges, li2021abstractive proposed an abstractive opinion tagging framework, named AOT-Net, that first predicts a salience score for each review, and given the salience scores, it groups all reviews into opinion clusters and ranks opinion clusters by cluster size. With the designed alignment feature and alignment loss, AOT-Net sequentially reads ranked opinion clusters and generates opinion tags with ranks. To generate opinion tags in a personalized way, zhao2022personaot selected the information that users are interested in from reviews and then generated a ranked list of aspect and opinion tag pairs. The authors tracked user preferences not only using explicit feedback, i.e., reviews, but also using implicit feedback such as clicks and purchases in a heterogeneous graph neural network model.
@@ 3.2 E-commerce users
  • Purchases. In e-commerce systems, purchases are very strong signals for recommendation. Most e-commerce platforms employ the Gross Merchandise Volume (GMV) as a gold standard for measuring success. GMV indicates the total amount of purchases from merchandise sales as the target of optimization of e-commerce [anderson2002new, lee2001visualization]. More and more recent studies utilize binary purchase information as the learning objective to characterize different levels of clicks [zhouwsdm2018].
@@ 3.2.1 From clicks to purchases
As an e-commerce user interacts with an item, they express a certain degree of interest in the item. When users browse an e-commerce platform, they may examine a specific item that is sufficiently relevant or intriguing. User clicks are an important signal for tracking a user’s interest [chuklin-click-2015]. A user’s online shopping behavior can be divided into two consecutive stages: item selection/clicks, and the decision to purchase the clicked item. Users may have different intentions while shopping online, e.g., some wish to make a purchase as soon as possible while others are just looking around so as to get inspired. Therefore, wu2018turning argued that clicks, as a kind of implicit feedback, should be integrated with other kinds of feedback to evaluate the “relevance” of items given a query on e-commerce portals.
@@
To boost sales, some online retailers modify the ranking of their items’ popularity with the usage of crowdsourcing platforms. For example, su2018detecting investigated and detected such kind of activities in e-commerce, e.g., crowd workers need to follow some particularly designed guidelines to disguise themselves as normal users. An example of this crowdsouring “add to favorites” task is shown in Fig. 3.16. By simultaneously manipulating a number of crowdsourcing tasks and collecting user behavior, the authors compare behavioral attributes between normal activities and spamming activities. Fig. 3.17 shows these comparisons in terms of four behavioral attributes: query length, page number, browse time (time period between search and click), and dwell time (on detailed item pages).
@@
Different recommendation scenarios on an e-commerce platforms may yield different types of user click and purchase behavior. E.g., clicks on the follow-up recommendation results after adding an item to the shopping cart, and clicks on the recommended results listed on an item’s detailed page [zhouwsdm2018]. The diversity in scenarios may make it harder to interpret clicks and their relation to purchase behavior. In e-commerce search and recommendation, a purchase action is a natural ground-truth label for a click. If a user ends up purchasing an item after clicking on it, such a click indicates the user’s strong interest in and satisfaction with the item. Accordingly, the conversion rate has been proposed as a important signal [zhouwsdm2018]:
@@
Purchase intent represents a predictive measure of subsequent purchasing behavior. Understanding purchase intent and how it is built up over time is important for personalized and contextualized e-commerce services. In recent years, many studies have explored the conversion from clicks to purchases [wen2018multi, guoziyi2018]. Moreover, to understand how user activities lead to purchase intent, both long and short-term purchase intent have been investigated [lo2016understanding, brown2003buying, kim2003combination, sismeiro2004modeling, swinyard2004activities, suh2004prediction, van2005predicting, young2004predicting].
@@
Studies focusing on short-term purchase intent analysis have investigated user demographics [young2004predicting], purchase patterns [kim2003combination], item attributes [brown2003buying, van2005predicting], and click streams [sismeiro2004modeling]. young2004predicting find that the transaction, cost, and incentive programs are important predictors for determining the short-term intention to purchase clothing, jewelry, and accessories on e-commerce portals. Furthermore, mcduff2015predicting presented a large-scale analysis of the connection between facial responses and purchases.
@@
lo2016understanding focused on long-term purchase intent analysis. The authors perform a large-scale long-term cross-platform study of user purchase intent and how it varies over time. They focus on four kinds of signals of user actions to detect purchase intent: closing-up on a piece of content, clicking through a link to an external website, searching for content, and saving content for later retrieval. The authors find that signals for purchase intent tend to slowly build up over time, and sharply increase about three to five days before a purchase. Moreover, users with a long-term purchase intent tend to save and click-through more content; these signals may be present for weeks before a purchase is made and they are amplified in the last three days before purchase.
@@
Social interactions can also be used to improve understanding of consumer behavior [guo2011role, gunawan2015viral, hajli2017social, testa2018social]. Users may consult their social network when they need to purchase something they are unfamiliar with. Thus, although social relations only provide implicit signals, they have been found to be useful to understand purchase decisions [guo2011role]. bhatt2010predicting find that purchase intent from highly connected individuals is correlated with adoption by users in their social circle. However, there is little evidence of social influence by these high degree individuals. The spread of purchase intent remains mostly local to first-adopters and their immediate friends. In a 2011 study of information passing in Taobao, guo2011role verified that implicit information passing is present in the network, and that communication between buyers is a fundamental driver of purchasing activity. zhang2013predicting presented a system to understand the relation between users’ Facebook profiles and purchase behaviors in eBay. Extensive analysis have been done on a benchmark dataset collected from Facebook and eBay; the authors find that there are significant correlations between social network information and online purchases.
@@ 3.2.2 User engagement and post-clicks
To illustrates post-click behavior on an e-commerce platform, Fig. 3.18 provides an example of observed data of a user during a short period. We see that the user first enters a product page for the iPhone 7 from a search result page. After reading the detailed description and comments, this user adds the item to their shopping cart. Then the user shifts to a page for the iPhone 6 from the search result page and reads the comments. After that, they browse a page devoted to iPhone 7 cases from the sales page and order the case. Finally, they jump to a page about Samsung Galaxy from the home page of the e-commerce site. During this period, two kinds of post-click behavior can be found: (i) from a coarse-grained perspective, the user interacted with the iPhone 7, the iPhone 6, iPhone 7 cases, and the Samsung Galaxy; and (ii) from a fine-grained perspective, each coarse-grained interaction includes a sequence of behavior that can indicate how the user locates the item page, whether the user clicks detailed information, whether a user adds-to-cart or orders an item, and how long the user dwells on an item [zhouwsdm2018].
@@
As mentioned in Section 3.2, typically, there are three browsing modules on e-commerce sites: the main page (including basic information, title, price, and pictures), the specification page (including more parameters and details), and the comment page. Fig. 3.19 illustrates the relations between clicks and these browsable components [zhouwsdm2018]. We see that a user is more likely to buy an item if they makes more clicks on its different browsable components, i.e., a user may gather basic information from the main item page, review feedback from the comment page, and click images to check if the item satisfies their requirements. guoziyi2018 showed how dwell time is related to clicks and browsable components; see Fig. 3.20. The dwell time on an item is related to how a user locates the item. As shown in Fig. 3.20(b), the longer the dwell time, the more likely a user would visit detailed components, including reading comments and specifications.
@@
zhouwsdm2018 investigated the relation between certain types of post-click behavior and the conversion rate (CVR). They find that the post-click behavior “Cart” has the highest conversion rate, which means if a user adds an item to the cart, they are more likely to order it in the end. Similarly, if a user enters an item page from the list of items in the cart, they are also very likely to order it. When the dwell time is outside a certain range, the conversion rate drops. If the user stays much longer than they need to finish the page, they might have transferred their attention offline. It is observed that users’ interactions often exhibit a monotonic structure, i.e., the presence of a more explicit interaction (such as reviews) necessarily implies the presence of a more implicit signal (such as clicks) [wan2018item].
@@ 3.3 Discussion
In this chapter, we have surveyed the infrastructure of e-commerce platforms, i.e., presentations and users. Specifically, we have introduced six information components that are widely applied in e-commerce platforms: search results, recommendation results, titles, product descriptions, question answering, and reviews. We have highlighted studies about user behavior in e-commerce, including user clicks, purchases, engagement, and post-click behavior.
@@
To gain a deeper understanding of information discovery on e-commerce platforms, we list three research questions to guide the following three chapters:
@@
We will address these questions through discussions in Chapter 4, Chapter 5, and Chapter 6, respectively. A considerable amount of relevant work about information components will be discussed in Chapter 7 as they have a clear connection to question answering and dialogue generation in e-commerce.
@@
In Chapter 3, we discussed work on e-commerce information infrastructures, focusing on e-commerce presentations, and on e-commerce users. The unique characteristics of e-commerce users make modeling for e-commerce users essential when attempting to understand and support information discovery [lo2016understanding, Huangwww2018]. E-commerce user modeling can be separated into two types: user behavior modeling and user profiling. Given specific user behavior in various scenarios, e.g., click behavior, purchasing behavior, and post-click behavior, user behavior modeling focuses on learning a model of user behavior to predict the user’s next preference. In contrast, user profiling aims to predict a user’s profile (e.g., age, gender, and occupation) given the user’s behavior records. In this chapter, we survey research on user behavior modeling and user profiling in e-commerce. First, we detail user behavior modeling approaches in Section 4.1. Next, we discuss studies on user profiling in e-commerce in Section 4.2. Lastly, Section 4.3 discusses emerging directions in e-commerce user modeling.
@@ 4.1.1 Click behavior modeling
Click actions recorded in query logs have successfully been applied to extract important features in the context of ranking scenarios [agichtein2006improving]. Regarding web search, click models have been proposed to help the search engine to accurately understand interactive user behavior [guo2009efficient, yilmaz2010expected, zhang2011user, chuklin-click-2015, borisov2016neural]. Early research on the topic aimed to track a user’s behavior by using probabilistic graphical models. More recently, neural networks have been applied to improve the performance of click models by representing user behavior to capture the user’s information needs [borisov2016neural]. Focusing on improving the effectiveness by exploiting information from user-system interactions, ferro2017including decided to explore embedding dynamic interactions into learning to rank frameworks. Thereafter, curriculum learning and continuation methods have been successfully applied to exploit user interactions and facilitate rank learning [ferro2019boosting].
Given the work mentioned above, click behavior modeling has received an increasing amount of attention in e-commerce scenarios [see, e.g., he2014practical, chapelle2015simple, chen2016deep, he2017neural, LiRCRLM17, wu2018turning, zhouwsdm2018, huang2019ecompred, gong2020edgerec, bian2021contra, wen2021hierarchy]. Viewing click prediction as a binary classification problem, the researchers who conducted those early studies employed logistic regression to predict whether an item will be clicked [richardson2007predicting], where handcrafted features are extracted from raw data to optimize a log-likelihood objective function for training. Latent factor optimization approaches, e.g., factorization machines [FM], have also been applied to utilize importance-aware and hierarchical structures purposed to manage dynamic user behavior [oentaryo2014predicting]. In Section 6.3.2, we detail studies about factorization machines in e-commerce recommendation.
@@
From shallow to deep models. CTR has been widely applied as an evaluation metric for click modeling in e-commerce portals. rendle2010pairwise proposed a tensor-based method for CTR prediction; Bayesian approaches have also been used effectively for CTR prediction [graepel2010web]. Starting in 2015, deep learning significantly improved CTR estimation by transferring traditional architectures and developing new ones. Deep neural networks effectively capture high-order feature interactions, resulting in better CTR prediction performance. zhang2016deep proposed a deep neural network to learn patterns from categorical feature interactions. Similarly, chen2016deep, zhu2017optimized employed neural network models with multiple fully-connected layers to predict user clicks. aryafar2017ensemble investigated CTR prediction in promoted listings by using an ensemble learning approach to leverage different signals of listings. Generally, these logistic regression models can effectively achieve memorization by applying cross-product transformations over sparse features. Recent work involves representing sparse features as dense vectors, which are concatenated to form an instance vector. This vector is then passed through a multi-layer perceptron, with a sigmoid output layer, to predict the click probability. These advancements have greatly enhanced model accuracy in CTR tasks [zhang2021deep].
Wide & Deep model. Modeling the interactions between features, especially the interactions between low-order and high-order features, is essential for click prediction. The Wide & Deep model [cheng2016wide] considers low- and high-order feature interactions simultaneously. Wide & Deep pursues the balance between memorization and generalization. Owing to its simple structure, the “strong” features (i.e., feature combinations) of Wide & Deep allow for the assignment of larger weights during training, thus endowing the model with stronger memory. Besides the deep component based on MLP, Wide & Deep consists of another components, the wide component. It is a generalized linear model with an input feature set that includes raw features and a feature that has been transformed by the cross-product transformation and is defined as
$\phi_{k}(x)=\prod_{i=1}^{d}x_{i}^{c_{ki}},\quad c_{ki}\in\{0,1\}$ (4.2)
where $c_{ki}$ is a Boolean variable that is 1 if the $i$-th feature is part of the $k$-th transformation $\phi_{k}$, and 0 otherwise. Such a transformation allows the model to capture the interactions between the binary features, and adds nonlinearity to the wide component. The overall model architecture of Wide & Deep is shown in Fig. 4.1. Wide & Deep has been verified as effective in e-commerce recommendation scenarios; more details are provided in Section 6.3.3.
@@
The Wide& Deep model is a representative of dual tower models for user behavior modeling. Similarly, DeepFM [guo2017deepfm], DCN [wang2017deep], xDeepFM [lian2018xdeepfm] and Autoint [song2020towards] are also proposed for CTR prediction. The deep neural network part in these dual tower models can always be regarded as a supplementary to learn the residual signal of the feature interaction layer to approach the label, which yields stable training and the improved performance. In contrast, Single-tower models like NFM [he2017neural] and Product-based Neural Network [Qu2018Product] have enhanced modeling capacity due to their more sophisticated network structures, which allow them to capture complex feature interactions. However, they often struggle with issues such as getting stuck in poor local minima and exhibit a heavy reliance on careful parameter initialization. This sensitivity to initialization can affect their training stability and convergence, making optimization more challenging compared to simpler models.
@@
Attention models for CTR. Attention neural networks have been proposed to enhance the performance of CTR prediction. The Deep Interest Network (DIN) [zhou2018deep] was the first model to introduce the attention network mechanism for user behavior modeling with CTR prediction. It assigns different weights to past behaviors based on their relevance to the target item. To capture dynamic interest evolution, the Deep Interest Evolution Network (DIEN) [zhou2019deep] was proposed, utilizing a two-layer GRU with an attentional update gate to model evolving user interests. Further advancements, like the Behavior Sequence Transformer and Deep Session Interest Network [feng2019deep], leverage self-attention to model behavior dependencies and session-based representations, showing the importance of attention mechanisms in CTR prediction [xiao2020deep]. Recent advances on user click model with attention models have focused on leveraging deep neural networks to capture complex interactions given user profiles, item attributes, and contextual features [hou2023deep]. These models have shown great potential in improving the accuracy and scalability.
@@
Memory-based models. With the accumulation of vast user behavior data on large e-commerce platforms, effectively handling long behavior sequences is increasingly important. However, many models such as DIN [zhou2018deep] still struggle with time complexity when processing such sequences. To address this, ren2019lifelong proposed the Hierarchical Periodic Memory Network, which uses a lifelong memory mechanism with multi-layer GRUs updating at different frequencies, capturing long-term and multi-scale temporal patterns. Similarly, the User Interest Center and Multi-channel User Interest Memory Network were designed to handle long-term user interest modeling, providing a more systematic, industrial-level approach [pi2019practice]. Recently, multi-interest networks have been studied to improve the robustness and consistency in user click modeling [cen2020controllable, chang2023twin]. To mitigate noisy correlations and user intent vanishing during this procedure, attribute transition graphs and matching among various patterns need to be constructed. To get this, liu2023enhancing characterized user intents with attribute patterns, where the frequent and compact attribute patterns are served as memory to augment session representations.
@@
Hybrid models combining multiple factors. To address the complexity of feature interactions, various hybrid models have been proposed. For example, the gradient boosting decision tree model [GBDT; chen2016xgboost] has been applied successfully to predict user clicks [he2014practical]. Fig. 4.2 illustrates the structure of a hybrid model with GBDT and logistic regression. The model concatenates the boosted decision trees, which transforms features and the sparse logistic regression classifier.
The input is a structured embedding $x=(e_{i_{1}},e_{i_{2}},\ldots,e_{i_{n}})$ for each item $x$, where $e_{i}$ refers to the $i$-th unit vector, and $i_{n}$ is the index of the categorical features. The output of the model is a binary label $y\in\{+1,-1\}$, which indicates a click or no click. Given a labeled pair $(x,y)$, the authors have denoted the linear combination of active weights as $s(y,x,w)$, which can be calculated as follows:

### v3

@@ 3.1.2.6 User reviews in e-commerce
Opinion tags refer to a ranked list of tags provided by the e-commerce platform that reflect the characteristics of reviews of an item; see Fig. 3.15. To assist consumers to quickly grasp a large number of reviews about an item, opinion tags are increasingly being applied by e-commerce platforms. Current mechanisms for generating opinion tags rely on either manual labelling or heuristic methods, which is time-consuming and ineffective. li2021abstractive introduce the abstractive opinion tagging task, where systems have to automatically generate a ranked list of opinion tags that are based on, but need not occur in, a given set of user-generated reviews.
The abstractive opinion tagging task comes with three main challenges: (i) the noisy nature of reviews; (ii) the formal nature of opinion tags vs. the colloquial language usage in reviews; and (iii) the need to distinguish between different items with very similar aspects. To address these challenges, li2021abstractive come up with an abstractive opinion tagging framework, named AOT-Net, that first predicts a salience score for each review, and given the salience scores, it groups all reviews into opinion clusters and ranks opinion clusters by cluster size. With the designed alignment feature and alignment loss, AOT-Net sequentially reads ranked opinion clusters and generates opinion tags with ranks. To generate opinion tags in a personalized way, zhao2022personaot select the information that users are interested in from reviews and then generated a ranked list of aspect and opinion tag pairs. The authors track user preferences not only using explicit feedback, i.e., reviews, but also using implicit feedback such as clicks and purchases in a heterogeneous graph neural network model.
@@ 3.2 E-commerce users
  • Purchases. In e-commerce systems, purchases are very strong signals for recommendation. Most e-commerce platforms employ the Gross Merchandise Volume (GMV) as a gold standard for measuring success. GMV indicates the total amount of purchases from merchandise sales as the target of optimization of e-commerce [anderson2002new, lee2001visualization]. Many recent studies use binary purchase information as the learning objective to characterize different levels of clicks [zhouwsdm2018].
@@ 3.2.1 From clicks to purchases
As an e-commerce user interacts with an item, they express a certain degree of interest in the item. When users browse an e-commerce platform, they may examine a specific item that is sufficiently relevant or intriguing. User clicks are an important signal for tracking a user’s interest [chuklin-click-2015]. A user’s online shopping behavior can be divided into two consecutive stages: item selection/clicks, and the decision to purchase the clicked item. Users may have different intentions while shopping online, e.g., some wish to make a purchase as soon as possible while others are just looking around so as to get inspired. Therefore, wu2018turning argue that clicks, as a kind of implicit feedback, should be integrated with other kinds of feedback to evaluate the “relevance” of items given a query on e-commerce portals.
@@
To boost sales, some online retailers modify the ranking of their items’ popularity with the usage of crowdsourcing platforms. For example, su2018detecting investigate and detect such kind of activities in e-commerce, e.g., crowd workers need to follow some particularly designed guidelines to disguise themselves as normal users. An example of this crowdsouring “add to favorites” task is shown in Fig. 3.16. By simultaneously manipulating a number of crowdsourcing tasks and collecting user behavior, the authors compare behavioral attributes between normal activities and spamming activities. Fig. 3.17 shows these comparisons in terms of four behavioral attributes: query length, page number, browse time (time period between search and click), and dwell time (on detailed item pages).
@@
Different recommendation scenarios on an e-commerce platforms may yield different types of user click and purchase behavior. E.g., clicks on the follow-up recommendation results after adding an item to the shopping cart, and clicks on the recommended results listed on an item’s detailed page [zhouwsdm2018]. The diversity in scenarios may make it harder to interpret clicks and their relation to purchase behavior. In e-commerce search and recommendation, a purchase action is a natural ground-truth label for a click. If a user ends up purchasing an item after clicking on it, such a click indicates the user’s strong interest in and satisfaction with the item. Accordingly, the conversion rate has been proposed as an important signal [zhouwsdm2018]:
@@
Purchase intent represents a predictive measure of subsequent purchasing behavior. Understanding purchase intent and how it is built up over time is important for personalized and contextualized e-commerce services. In recent years, many studies have explored the conversion from clicks to purchases [wen2019multi, guoziyi2018]. Moreover, to understand how user activities lead to purchase intent, both long and short-term purchase intent have been investigated [lo2016understanding, brown2003buying, kim2003combination, sismeiro2004modeling, swinyard2004activities, suh2004prediction, van2005predicting, young2004predicting].
@@
Studies focusing on short-term purchase intent analysis have investigated user demographics [young2004predicting], purchase patterns [kim2003combination], item attributes [brown2003buying, van2005predicting], and click streams [sismeiro2004modeling]. young2004predicting find that the transaction, cost, and incentive programs are important predictors for determining the short-term intention to purchase clothing, jewelry, and accessories on e-commerce portals. Furthermore, mcduff2015predicting present a large-scale analysis of the connection between facial responses and purchases.
@@
lo2016understanding focus on long-term purchase intent analysis. The authors perform a large-scale long-term cross-platform study of user purchase intent and how it varies over time. They focus on four kinds of signals of user actions to detect purchase intent: closing-up on a piece of content, clicking through a link to an external website, searching for content, and saving content for later retrieval. The authors find that signals for purchase intent tend to slowly build up over time, and sharply increase about three to five days before a purchase. Moreover, users with a long-term purchase intent tend to save and click-through more content; these signals may be present for weeks before a purchase is made and they are amplified in the last three days before purchase.
@@
Social interactions can also be used to improve understanding of consumer behavior [guo2011role, gunawan2015viral, hajli2017social, testa2018social]. Users may consult their social network when they need to purchase something they are unfamiliar with. Thus, although social relations only provide implicit signals, they have been found to be useful to understand purchase decisions [guo2011role]. bhatt2010predicting find that purchase intent from highly connected individuals is correlated with adoption by users in their social circle. However, there is little evidence of social influence by these high degree individuals. The spread of purchase intent remains mostly local to first-adopters and their immediate friends. In a 2011 study of information passing in Taobao, guo2011role verify that implicit information passing is present in the network, and that communication between buyers is a fundamental driver of purchasing activity. zhang2013predicting present a system to understand the relation between users’ Facebook profiles and purchase behaviors in eBay. Extensive analyses have been done on a benchmark dataset collected from Facebook and eBay; the authors find that there are significant correlations between social network information and online purchases.
@@ 3.2.2 User engagement and post-clicks
To illustrates post-click behavior on an e-commerce platform, Fig. 3.18 provides an example of observed data of a user during a short period. We see that the user first enters a product page for the iPhone 7 from a search result page. After reading the detailed description and comments, this user adds the item to their shopping cart. Then, the user shifts to a page for the iPhone 6 from the search result page and reads the comments. After that, they browse a page devoted to iPhone 7 cases from the sales page and order the case. Finally, they jump to a page about the Samsung Galaxy from the home page of the e-commerce site. During this period, two kinds of post-click behavior can be found: (i) from a coarse-grained perspective, the user interacted with the iPhone 7, the iPhone 6, iPhone 7 cases, and the Samsung Galaxy; and (ii) from a fine-grained perspective, each coarse-grained interaction includes a sequence of behavior that can indicate how the user locates the item page, whether the user clicks detailed information, whether a user adds-to-cart or orders an item, and how long the user dwells on an item [zhouwsdm2018].
@@
As mentioned in Section 3.2, typically, there are three browsing modules on e-commerce sites: the main page (including basic information, title, price, and pictures), the specification page (including more parameters and details), and the comment page. Fig. 3.19 illustrates the relations between clicks and these browsable components [zhouwsdm2018]. We see that a user is more likely to buy an item if they produce more clicks on its different browsable components, i.e., a user may gather basic information from the main item page, review feedback from the comment page, and click images to check if the item satisfies their requirements. guoziyi2018 show how dwell time is related to clicks and browsable components; see Fig. 3.20. The dwell time on an item is related to how a user locates the item. As shown in Fig. 3.20(b), the longer the dwell time, the more likely a user would visit detailed components, including reading comments and specifications.
@@
zhouwsdm2018 investigate the relation between certain types of post-click behavior and the conversion rate (CVR). They find that the post-click behavior “Cart” has the highest conversion rate, which means if a user adds an item to the cart, they are more likely to order it in the end. Similarly, if a user enters an item page from the list of items in the cart, they are also very likely to order it. When the dwell time is outside a certain range, the conversion rate drops. If the user stays much longer than they need to finish the page, they might have transferred their attention offline. It is observed that users’ interactions often exhibit a monotonic structure, i.e., the presence of a more explicit interaction (such as reviews) necessarily implies the presence of a more implicit signal (such as clicks) [wan2018item].
@@ 3.3 Discussion
In this section, we have surveyed the infrastructure of e-commerce platforms, i.e., presentations and users. Specifically, we have introduced six information components that are widely applied on e-commerce platforms: search results, recommendation results, titles, product descriptions, question answering, and reviews. We have highlighted studies about user behavior in e-commerce, including user clicks, purchases, engagement, and post-click behavior.
@@
To gain a deeper understanding of information discovery on e-commerce platforms, we list three research questions to guide the following three sections:
@@
We will address these questions through discussions in Section 45, and 6, respectively. A considerable amount of relevant work about information components will be discussed in Section 7 as they have a clear connection to question answering and dialogue generation in e-commerce.
@@
In Section 3, we discussed work on e-commerce information infrastructures, focusing on e-commerce presentations, and on e-commerce users. The unique characteristics of e-commerce users make modeling for e-commerce users essential when attempting to understand and support information discovery [lo2016understanding, Huangwww2018]. E-commerce user modeling can be separated into two types: user behavior modeling and user profiling. Given specific user behavior in various scenarios, e.g., click behavior, purchasing behavior, and post-click behavior, user behavior modeling focuses on learning a model of user behavior to predict the user’s next preference. In contrast, user profiling aims to predict a user’s profile (e.g., age, gender, and occupation) given the user’s behavior records. In this section, we survey research on user behavior modeling and user profiling in e-commerce. First, we detail user behavior modeling approaches in Section 4.1. Next, we discuss studies on user profiling in e-commerce in Section 4.2. Lastly, Section 4.3 discusses emerging directions in e-commerce user modeling.
@@ 4.1.1 Click behavior modeling
Click actions recorded in query logs have successfully been applied to extract important features in the context of ranking scenarios [agichtein2006improving]. Regarding web search, click models have been proposed to help the search engine understand interactive user behavior [guo2009efficient, yilmaz2010expected, zhang2011user, chuklin-click-2015, borisov2016neural]. Early research on the topic aimed to track a user’s behavior by using probabilistic graphical models. More recently, neural networks have been applied to improve the performance of click models by representing user behavior to capture the user’s information needs [borisov2016neural]. Focusing on improving the effectiveness by exploiting information from user-system interactions, ferro2017including explore embedding dynamic interactions into learning to rank frameworks. Thereafter, curriculum learning and continuation methods have been successfully applied to exploit user interactions and facilitate rank learning [ferro2019boosting].
Given the work mentioned above, click behavior modeling has received an increasing amount of attention in e-commerce scenarios [see, e.g., he2014practical, chapelle2015simple, chen2016deep, he2017neural, LiRCRLM17, wu2018turning, zhouwsdm2018, huang2019ecompred, gong2020edgerec, bian2021contra, wen2021hierarchy]. Viewing click prediction as a binary classification problem, the researchers who conducted those early studies employed logistic regression to predict whether an item will be clicked [richardson2007predicting], where handcrafted features are extracted from raw data to optimize a log-likelihood objective function for training. Latent factor optimization approaches, e.g., factorization machines [FM], have also been applied to use importance-aware and hierarchical structures purposed to manage dynamic user behavior [oentaryo2014predicting]. In Section 6.3.2, we detail studies about factorization machines in e-commerce recommendation.
@@
From shallow to deep models. CTR has been widely applied as an evaluation metric for click modeling in e-commerce portals. rendle2010pairwise introduce a tensor-based method for CTR prediction; Bayesian approaches have also been used effectively for CTR prediction [graepel2010web]. Starting in 2015, deep learning significantly improved CTR estimation by transferring traditional architectures and developing new ones. Deep neural networks effectively capture high-order feature interactions, resulting in better CTR prediction performance. zhang2016deep describe a deep neural network to learn patterns from categorical feature interactions. Similarly, chen2016deep, zhu2017optimized employ neural network models with multiple fully-connected layers to predict user clicks. aryafar2017ensemble investigate CTR prediction in promoted listings by using an ensemble learning approach to use different signals of listings. Generally, these logistic regression models can effectively achieve memorization by applying cross-product transformations over sparse features. More recent work involves representing sparse features as dense vectors, which are concatenated to form an instance vector. This vector is then passed through a multi-layer perceptron, with a sigmoid output layer, to predict the click probability. These advancements have greatly enhanced model accuracy in CTR tasks [zhang2021deep].
Wide & Deep model. Modeling the interactions between features, especially the interactions between low-order and high-order features, is essential for click prediction. The Wide & Deep model [cheng2016wide] considers low- and high-order feature interactions simultaneously. Wide & Deep pursues the balance between memorization and generalization. Owing to its simple structure, the “strong” features (i.e., feature combinations) of Wide & Deep allow for the assignment of larger weights during training, thus endowing the model with stronger memory. Besides the deep component based on an MLP, Wide & Deep consists of another component, the wide component. It is a generalized linear model with an input feature set that includes raw features and a feature that has been transformed by the cross-product transformation and is defined as
$\phi_{k}(x)=\prod_{i=1}^{d}x_{i}^{c_{ki}},\quad c_{ki}\in\{0,1\},$ (4.2)
where $c_{ki}$ is a Boolean variable that is 1 if the $i$-th feature is part of the $k$-th transformation $\phi_{k}$, and 0 otherwise. Such a transformation allows the model to capture the interactions between the binary features, and adds nonlinearity to the wide component. The overall model architecture of Wide & Deep is shown in Fig. 4.1. Wide & Deep has been shown to be effective in e-commerce recommendation scenarios; more details are provided in Section 6.3.3.
@@
The Wide& Deep model is a representative of dual tower models for user behavior modeling. Similarly, DeepFM [guo2017deepfm], DCN [wang2017deep], xDeepFM [lian2018xdeepfm] and Autoint [song2020towards] have also been put forward for CTR prediction. The deep neural network part in these dual tower models can always be regarded as a supplementary to learn the residual signal of the feature interaction layer to approach the label, which yields stable training and the improved performance. In contrast, single-tower models like NFM [he2017neural] and the product-based neural network [Qu2018Product] have enhanced their modeling capacity due to their more sophisticated network structures, which allow them to capture complex feature interactions. However, they often struggle with issues such as getting stuck in poor local minima and exhibit a heavy reliance on careful parameter initialization. This sensitivity to initialization can affect their training stability and convergence, making optimization more challenging than for simpler models.
@@
Attention models for CTR. Attention neural networks have been proposed to enhance the performance of CTR prediction. The deep interest network (DIN) [zhou2018deep] is the first model to introduce the attention network mechanism for user behavior modeling with CTR prediction. It assigns different weights to past behaviors based on their relevance to the target item. To capture dynamic interest evolution, the deep interest evolution network (DIEN) [zhou2019deep] has been proposed; it uses a two-layer GRU with an attentional update gate to model evolving user interests. Further advancements, like the behavior sequence transformer and the deep session interest network [feng2019deep], use self-attention to model behavior dependencies and session-based representations, showing the importance of attention mechanisms in CTR prediction [xiao2020deep]. More recent advances in user click models with attention have focused on using deep neural networks to capture complex interactions given user profiles, item attributes, and contextual features [hou2023deep]. These models have shown great potential in improving the accuracy and scalability.
@@
Memory-based models. With the accumulation of large amounts of user behavior data on large e-commerce platforms, effectively handling long behavior sequences is increasingly important. However, many models such as DIN [zhou2018deep] struggle with the time complexity when processing such sequences. To address this, ren2019lifelong introduce the hierarchical periodic memory network; it uses a lifelong memory mechanism with multi-layer GRUs updating at different frequencies, capturing long-term and multi-scale temporal patterns. Similarly, the user interest center and the multi-channel user interest memory network are designed to handle long-term user interest modeling, providing a more systematic, industrial-level approach [pi2019practice]. Multi-interest networks have also been studied to improve the robustness and consistency in user click modeling [cen2020controllable, chang2023twin]. To mitigate noisy correlations and user intent vanishing during this procedure, attribute transition graphs and matching among various patterns need to be constructed. To this end, liu2023enhancing characterize user intents with attribute patterns, where the frequent and compact attribute patterns serve as memory to augment session representations.
@@
Hybrid models combining multiple factors. To address the complexity of feature interactions, various hybrid models have been proposed. For example, the gradient boosting decision tree model [GBDT; chen2016xgboost] has been applied successfully to predict user clicks [he2014practical]. Fig. 4.2 illustrates the structure of a hybrid model with GBDT and logistic regression. The model concatenates the boosted decision trees, which transforms features and the sparse logistic regression classifier. The input is a structured embedding $x=(e_{i_{1}},e_{i_{2}},\ldots,e_{i_{n}})$ for each item $x$, where $e_{i}$ refers to the $i$-th unit vector, and $i_{n}$ is the index of the categorical features. The output of the model is a binary label $y\in\{+1,-1\}$, which indicates a click or no click. Given a labeled pair $(x,y)$, the authors denote the linear combination of active weights as $s(y,x,w)$, which can be calculated as follows: